OpenStudy (anonymous):

Solve: 2(x + 2) − (x + 1) = 8

OpenStudy (chihiroasleaf):

first, use distributive property 2(x+2)-(x+1)=8 2x+2.2 -x-1=8 2x+4-x-1=8 next, simplify.., what do you get?

OpenStudy (anonymous):

x+4-1=8

OpenStudy (chihiroasleaf):

yes.., can you simplify it again?

OpenStudy (anonymous):

x+3=8 x=5

OpenStudy (anonymous):

Thank you

OpenStudy (chihiroasleaf):

well done... you're welcome :)